Suds & Cuts is a local pet grooming shop owned by Collin Bark. Collin has prepared the following standard cost card for each dog bath given: Standard Quantity Standard Rate Standard Unit Cost Shampoo 1.7 oz. $ 0.70 per oz. $ 1.19 Water 15 gal. $ 0.05 per gal. 0.75 Direct labor 1.05 hr. $ 10.00 per hr. 10.50 During the month of July, Collin’s employees gave 350 baths. The actual results were 725 ounces of shampoo used (cost of $558.25), 6,500 gallons of water used (cost of $455), and labor costs for 290 hours (cost of $3,306).

Required: 1. Calculate Suds & Cuts direct materials variances for both shampoo and water for the month of July. (Round your answers to 2 decimal places. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable.)

Direct Materials Per Variance? Direct materials quantity variance?
Answer #1

The formulas for calcuating direct material price and direct material quantity variance are given below:

Direct Material Price Variance = Actual Material*(Actual Rate - Standard Rate)

Direct Material Quantity Variance = Standard Rate*(Actual Quantity - Standard Quantity for Actual Baths)

___________

Shampoo:

Direct Material Price Variance = 725*(558.25/725 - .70) = $50.75 (U)

Direct Material Quantity Variance = .70*(725 - 350*1.70) = 91 (U)

___________

Water:

Direct Material Price Variance = 6,500*(455/6,500 - .05) = 130 (U)

Direct Material Quantity Variance = .05*(6,500 - 350*15) = 62.50 (U)